Complete the factorization of 36⁢x^2+12⁢x+1 by supplying the missing value.

36⁢x^2+12⁢x+1=(?x+1)2

The solution is
Answer #1

36x²+12x+1 = (6x)²+ 2×6x×1 +1²

[Splitting it into form of (a²+b²+2ab) = (a+b)²]

= (6x+1)²

Therefore, the missing value is 6.